A is a solution containing 0.0905 moldm3 of NHO3 B was prepared by diluting 50.0cm3 of a saturation of Ma2Co2 at room temperature to 1000 cm-3 25cm3 portion of B was found to required and average 24.90cm3 of A for complete neutralization using methl orange as the indicator, the data above. Determine
1. Amount of HNO3 in average volume of A used
2. Concentration of B in moldm3
3. Solubility of Na2Co3 in moldm3
4. Molar mass of Na2Co3 that will obtain evaporating 1dm3 of the solution to dryness by evaporating 2HNO3+NaCo _ 2NaNo3
1
Expert's answer
2021-09-20T02:27:08-0400



Average volume = "\frac{25+24.90}{2}=24.9cm^3"


Moles of A = "\frac{0.0905×1000}{1000}=0.0905moles"


Molarity of B ="0.0905×\frac{1000}{24.9}=3.63Moldm^3"